Hi there. I need some help.
solve for x or y.
2x-3y=2
7x-7y=5
I dont understand the -2

Answer by PRMath(133)

solve for x or y.
2x-3y=2
7x-7y=5

I'm not sure if you want each equation solved for x or y, or whether this is a system. I think this is a system, so if it isn't, just re-submit your question and I'm sure you'll get an answer.
There are a few ways you can solve this but I will show you how to solve this system by elimination.........

2x - 3y = 2 (What if you multiply this whole line by 7?)
7x - 7y = 5 (What if you multiply this whole line by -2?)

Here are your new equations after multiplying by 7 (top line) and -2 (bottom line)

14x - 21y = 14
-14x + 14y = -10 (now add these two equations)
________________________
0x - 7y = 4
y = -4/7

Now you can "plug" -4/7 into one of your original equations. I'll choose the first one:

2x-3y=2 (original equation)
2x - 3(-4/7) = 2 (plug in -4/7)
2x + 12/7 = 2 (multiply -3 times -4/7)
2x = 2 - 12/7 (subtract 12/7 from both sides)
2x = 2/1 - 12/7 (showing that 2 is the same as 2/1)
2x = 14/7 - 12/7 (finding a common denominator for combining fractions)
2x = 2/7 (subtraction)
x = 2/7 divided by 2 (divide by 2 on both sides)
x = 2/7 times 1/2 (when we divide fractions, we multiply by the reciprocal)
x = 2/14 = 1/7

Now if your question was an equation of a line, that's something else...
I hope this is helpful....